OpenStudy (anonymous):

Jake has a bag of 50 beads, of which some are red and the remaining are yellow. Jake randomly pulls out a bead from the bag, records the color, and replaces it in the bag. Jake has already recorded 6 red and 9 yellow beads. Based on the result, what is most likely the number of red beads in the bag? 10 20 30 40

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

10

OpenStudy (anonymous):

A vendor has 12 balloons for sale: 6 are yellow, 2 are red, and 4 are green. A balloon is selected at random and sold. If the balloon sold is yellow, what is the probability that the next balloon selected at random is also yellow? 6 over 12 5 over 12 5 over 11 6 over 11

OpenStudy (anonymous):

5 over 11

OpenStudy (anonymous):

The figure below shows a shaded rectangular region inside a large rectangle: A rectangle of length 10 units and width 5 units is shown. Inside this rectangle is another rectangle of length 6 units and width 4 units placed symmetrically inside the larger rectangle. The smaller rectangle is shaded gray. What is the probability that a point chosen inside the large rectangle is not in the shaded region? 24% 48% 52% 64%

OpenStudy (anonymous):

Thank you for helping me.

OpenStudy (anonymous):

\[\frac{50-24}{50}=\frac{13}{25}=0.52=52 \% \]